[Hallc_running] [New Logentry] RC Update

latif at jlab.org latif at jlab.org
Fri Nov 9 11:45:02 EST 2018

Logentry Text:
- We finished kinematic group 2 data early this morning.

- Some slow data taking was attributed to excessive SHMS HG Cerenkov cut in the bean counter. (See https://logbooks.jlab.org/entry/3623448) 

- The SHMS polarity was changed and the momentum was set to -4.262 GeV/c to take data for kinematic group 1. (see https://logbooks.jlab.org/entry/3623454)

- During the SHMS polarity change, Jay was notified (as per his previous suggestion) to check BPM. They did a quick test and made a logbook entry about it. (See https://logbooks.jlab.org/entry/3623453)

- The MCC has swapped the plan for Monday and Tuesday. Now they are planning to do the system recovery on Monday (~ 4 hours) and extended beam studies (~ 8 hours) on Tuesday. They expect halls A, B and C to stay locked up during the tests.

- During the beam studies, MCC will send around 85uA to Hall-C which would require the carbon hole or empty target.

- Currently, we are taking data for kinematic setting 1-3.

We still need owl shift takers for Saturday, Monday and Tuesday.

The daily run plan is posted here: https://hallcweb.jlab.org/wiki/index.php/RC_daily_meetings_and_daily_runplan_fall_2018

We will meet today at 4:30 PM in the counting house (second floor) to discuss the update and run plan.
Bluejeans link: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3623516
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181109/061fdcfc/attachment.html>

More information about the Hallc_running mailing list